How many took Math as their only subject? English only? Filipino only

1
Expert's answer
2021-11-09T12:48:23-0500

Incomplete question:

Let us take an example related to this problem:

A survey on subjects being taken by 250 college students in La Union revealed the following information: (Subject-number of students taking the subject) Mathematics - 145; Filipino - 90; English - 88; Math and Filipino - 25; Filipino and English - 38; Math and English - 59; Math, Filipino and English 15.

Solution:


Venn diagram:



Number of students who has taken Math as their only subject "= 145 -(10+15+44)=76"

Number of students who has taken English as their only subject "= 88 -(44+15+23)=6"

Number of students who has taken Filipino as their only subject "= 90 -(10+15+23)=42"